When it was first introduced as an antidepressant ten years ago, the average patient with a Prozac prescription was 32 years old. In a recent random sample of 20 prescription holders, the average age was 25 with a standard deviation s = 10 years. Use this information to answer the following.

11. If you wanted to determine if Prozac users are younger now on average, which of the following would be your null hypothesis?

A.☐= 25 B.☐= 32. C. ☐= 7. D.☐< 32 E.☐1-2 = 0

12. Which would be the logical alternative hypothesis? A.☐> 25
B.☐1- 2<0

C.☐D =7 D.☐< 32

E.☐32

13. If you assumed that the population of Prozac users had a normally distributed age distribution (because the sample is small), which of the following would be your critical value for determining whether to reject or not reject the null using alpha = .10

A. ☐ 1.330 B. ☐ -1.729 C. ☐ -1.325 D. ☐ 1.328 E. ☐ -1.328

14. The t score you would calculate from your sample data would be A. ☐2.236

B.☐.70 C.☐.35. D. ☐ -2.236. E. ☐ -3.13
Answer #1

Given:

= 32, n = 20, = 25, SD(S) = 10, = 0.10

11) Null Hypothesis:

ANSWER: B

Ho: = 32

12) Alternative Hypothesis:

ANSWER: D

Ha: < 32

13)

Degree of Freedom = n-1 = 20-1 = 19

Critical value: = 0.10

...................Using t table

ANSWER: B -1.729

14)

Test statistic:

ANSWER: E -3.13